If you use a marble of radius 13 mm for mercury, how far from the sun must you place your outpost planet?


1
Expert's answer
2021-06-13T07:09:02-0400

Solution.

"r=13mm=0.013m;"

"R_M=2440km=2440000m;"

"D=57.9\sdot 10^9m;"

"\dfrac{R_M}{r}=\dfrac{D}{d}\implies d=\dfrac{D\sdot r}{R_M};"

"d=\dfrac{57.9\sdot10^9m\sdot0.013m}{2440000m}=308m;"

Answer: "d=308m."